# New Advances in Phonons: From Band Topology to Quasiparticle Chirality

### v1

@@ II.2.3 Unconventional Dirac Phonons
As discussed in the previous subsection, surface arcs arise when a pair of Weyl phonons with opposite Chern numbers project to distinct momenta in the surface Brillouin zone (BZ), as illustrated in Figs. 7 (a1)-(a3). Conversely, if two Weyl points with opposite chirality project to the same momentum, the surface arcs vanish, as shown in Fig. 7 (a4). When such a pair of Weyl phonons carrying opposite Chern numbers approaches each other under $\mathcal{PT}$ symmetry, they coalesce into a Dirac node accompanied by two counter-propagating helicoidal surface states. Crucially, this $\mathcal{PT}$-protected Dirac phonon carries a net monopole charge of zero, rendering it topologically trivial and devoid of protected surface states in Dirac semimetals Kargarian et al. (2016); Le et al. (2018). Introducing a symmetry-preserving perturbation hybridizes the anti-parallel surface states, lifting their degeneracy and generating gapped, topologically trivial surface states Kargarian et al. (2016); Le et al. (2018). Despite this straightforward theoretical picture, two critical challenges remain unresolved before the year of 2018:
@@
(p1) In electronic systems, double Fermi arcs have predominantly been proposed in Dirac semimetals like Na3Bi Wang et al. (2012); Xiong et al. (2015); Liu et al. (2014b) and Cd3As2Wang et al. (2013); Liu et al. (2014a); Neupane et al. (2014); Yi et al. (2014); He et al. (2014); Borisenko et al. (2014); Li et al. (2016), they serve as critical experimental signatures of their topological nature. These arcs arise from the hybridization of two anti-parallel surface states, which remains incomplete due to their weak hybridization. A key unresolved question is identifying the conditions that lead to fully gapped surface states (and the absence of Fermi arcs). In other words, what is the microscopic mechanism for obtaining a strong hybridization of two anti-parallel surface states in Dirac phonon systems?
@@
To address (p1), Le et al. (2018) pointed out that the deformation of the anti-parallel surface states (termed double Fermi arcs in their study) arises due to a significant $k^{3}$ term that respects all crystal symmetries. They illustrated this with first-principle calculations on $\beta$-CuI, for which the effective Hamiltonian at the Dirac point can be formulated as $H_{\text{eff}}(\bm{k})$:
@@
To address (p2), a new type of topological invariant termed $\mathbb{Z}_{2}$-type monopole charge $Q$ must be introduced for topological Dirac phonons, since the $\mathbb{Z}$-type monopole charge $C$ is zero. Fang et al. (2016); Zhang et al. (2022b); Zhang and Murakami (2023); Yukitake et al. (2025b) have highlighted that in systems exhibiting $\mathcal{TG}$ symmetry, where $\mathcal{G}$ signifies the glide mirror symmetry, this $\mathbb{Z}_{2}$ monopole charge $Q$ can be defined Fang et al. (2016); Cheng et al. (2020b); Zhang et al. (2022b); Zhang and Murakami (2023); Yukitake et al. (2025b); Su et al. (2022); Cai et al. (2020); Qian et al. (2023); Lu et al. (2016a); Kim et al. (2021); Yukitake et al. (2025a). When $Q\neq 0$, double helicoid surface states emerge, as shown in Fig. 10 (a3). Conversely, if $Q=0$, the Dirac phonon is trivial with gapped surface states. Details are as follows.
@@
Figure 9: The $\mathbb{Z}_{2}$ monopole charge $Q$ is defined using specific surfaces and lines. In (a), $Q$ is defined on a surface $S$ that does not intersect with the plane $k_{z}=0$. This surface $S$ is closed under $\hat{\Theta}$, with the half part $S_{+}$ ($k_{z}\geq\pi$) shown in (b). The boundary of $S_{+}$, denoted as $\partial S_{+}=l_{1}\cup l_{2}$, consists of two lines. Examples of such surfaces $S$ are provided in (c) and (d), where $S$ is defined by a torus $T$ and a sphere $S^{2}$, respectively. (e) illustrates a closed line $l$ used to define the quantity $P[l]$. Adapted from Ref. Yukitake et al. (2025b).
@@
Yukitake et al. (2025b) considered a system that possess a single glide symmetry operation, denoted as $\mathcal{G}_{y}=\{M_{y}|00\frac{1}{2}\}$, as shown in Fig. 10 (a). When the time-reversal symmetry, $\mathcal{T}$, is a symmetry of the system, the operation $\tilde{\Theta}_{y}$ =${\mathcal{T}}$$\mathcal{G}_{y}$ is also conserved. It is found that $\tilde{\Theta}_{y}^{2}=e^{-ik_{z}}$, indicating that the $k_{z}=\pi$ ($-\pi$) plane is distinctive, as it satisfies the condition $\tilde{\Theta}_{y}^{2}=-1$. In this context, Yukitake et al. (2025b) examined the $\mathbb{Z}_{2}$-type Dirac points on the $k_{z}=\pi$ plane in the 3D BZ. These two Dirac points are connected by ${\mathcal{T}}$ symmetry, precluding their occurrence at time-reversal-invariant momenta (TRIM). Consequently, $\mathbb{Z}_{2}$ Dirac points typically emerge along high-symmetry lines in systems possessing a single glide mirror symmetry, and at non-TRIM high-symmetry points in systems with two vertical glide mirror symmetries Fang et al. (2016); Zhang et al. (2022b); Zhang and Murakami (2023); Yukitake et al. (2025b); Kim et al. (2021); Yukitake et al. (2025a). Subsequently, we will review the convention-independent $\mathbb{Z}_{2}$ monopole charge $Q[S]$ associated with an oriented closed surface $S$, which adheres to the following criteria as established by Yukitake et al. (2025b): (i) The surface $S$ is closed under the operation $\hat{\Theta}:(k_{x},k_{y},k_{z})\mapsto(-k_{x},k_{y},-k_{z})$, (ii) The surface $S$ does not intersect with the plane defined by $k_{z}=0$ (iii) The system exhibits a bulk energy gap over the entire surface $S$.
@@
Analogous to $\mathbb{Z}$-type monopole charge ${C}$, the $\mathbb{Z}_{2}$-type monopole charge ${Q}$ is defined in terms of wavefunctions over a sphere $S$ that encloses the Dirac point, as described in Fang et al. (2016). Thus, ${Q}$ can be expressed as:
[…]
@@
$(-1)^{Q}=\prod_{l\in\partial S_{n}}\frac{\operatorname{Pf}{\omega}[l(0)]}{\sqrt{\det\omega[l(0)]}}\frac{\operatorname{Pf}{\omega}[l(\pi)]}{\sqrt{\det\omega[l(\pi)]}}.$ (38)
@@
Thus, $Q[T]$ is the difference of $\tilde{\Theta}$-polarization, and it serves to establish the bulk-surface correspondence of $\mathbb{Z}_{2}$ Dirac points by considering the surface $T$ as a 2D BZ $[-\pi,\pi]\times[-\pi,\pi]$ in the $(s,t)$ parameter space, as shown in Fig. 10 (a1). Given that $\tilde{\Theta}^{2}=-1$ along the lines where $t=0$ and $t=\pi$, the anti-unitary operator $\tilde{\Theta}$ behaves analogously to time-reversal symmetry in $\mathcal{T}$-symmetric $\mathbb{Z}_{2}$ topological insulators Fu and Kane (2007). This implies that the energy bands, both from the bulk and the surface, will exhibit Kramers’s degeneracy at the points at $t=0$ and $t=\pi$. Upon projection onto the (100) surface BZ, the torus $T$ is projected onto a circle path $c$, as marked by the orange circle in Fig. 10 (a1). If the $\mathbb{Z}_{2}$ charge $Q[T]$ is nontrivial, an odd number of $\mathbb{Z}_{2}$ Dirac points are enclosed within the torus. Consequently, a pair of helical surface states will emerge on the circle $C$, as shown by the red and blue bands in Fig. 10 (a2). These two surface states are degenerate at $t=0$, resulting in a degenerate line along the boundary of the surface BZ. Adjusting the radius of circle $C$ allows for the emergence of two helical surface states, as shonw in Fig. 10 (a3). The intersection of these two surface states is protected by the symmetry $\mathcal{GT}$, leading to the formation of double helical (anti-parallel) surface states. Figure 10 (a4) illustrates the Fermi surface of a system hosting two $\mathbb{Z}_{2}$ Dirac points, featuring two Fermi arcs that link the projection of the Dirac points.
@@
$\mathbb{Z}_{2}$-type Dirac points may reside at various momenta, subject only to the constraint of satisfying the $\tilde{\Theta}_{y}$ symmetry, which is associated with a high-symmetry line in the BZ. When two Dirac points are projected onto distinct momenta within the surface BZ, a pair of anti-parallel helical surface states emerges. These states are topologically protected by the charge ${Q}$, as shown in Fig. 10 (b2). When two $\tilde{\Theta}_{i,j}$ symmetries are present, with $i,j\in\{x,y,z\}$, Dirac points are confined to the high-symmetry points within the BZ, as listed in Tab. 10. In this case, if two Dirac points are projected onto the identical momentum within the surface BZ, the quadruple surface states appear, as shown in Fig. 10 (b3).
@@
In certain instances, a $\mathbb{Z}_{2}$ Dirac point may split into a pair of Weyl dipoles or transform into a $\mathbb{Z}_{2}$ nodal ring, a transformation that depends on the crystalline symmetries inherent to the system. All of the topological band crossings mentioned here can be cataloged by the $\mathbb{Z}_{2}$-type monopole charge ${Q}$, associated with topologically protected surface states Fang et al. (2016); Cheng et al. (2020b); Zhang et al. (2022b); Zhang and Murakami (2023); Yukitake et al. (2025b); Su et al. (2022); Cai et al. (2020); Qian et al. (2023); Kim et al. (2021); Yukitake et al. (2025a); Lu et al. (2016a).
@@
Figure 10: “Bulk-surface correspondence” for the $\mathbb{Z}_{2}$ monopole charge $Q[T]$, defined on the torus $T$. (a1) Bulk and (100) surface Brillouin Zones. A $\mathbb{Z}_{2}$ Dirac point is situated within the $k_{z}=\pi$ plane of the torus $T$ and is projected onto the circle $\mathcal{C}$. The $k_{z}=\pi$ plane is characterized by $\mathcal{TG}^{2}=-1$. The torus may be conceptualized as a 2D rillouin Zone in the parameter space of ($s,t$), where Krammers’ degeneracy appears at $t=\pi,0$. (a2) When a $\mathbb{Z}_{2}$ Dirac point is enclosed within $T$, the surface states along the loop $\mathcal{C}$ exhibit helical characteristics and are degenerate at $t=0$ ($t=\pi$). (a3) By altering the radius of the loop $\mathcal{C}$, double-helicolid surface states is obtained, associated with two Fermi arcs shown in (a4). The charge $Q[T]$ is applicable to systems with $\mathbb{Z}_{2}$-charged nodal ring or a Weyl dipole, which can interconvert with Dirac points. (b1) Represents the topological surface state connecting a pair of Weyl phonons with opposite chirality. (b2) and (b3) are the double and quadruple helical surface states, respectively, which are attributed to the $\mathbb{Z}_{2}$ Dirac phonons. (a) is adapted from Ref. Yukitake et al. (2025b) and (b) is adapted from Ref. Zhang and Murakami (2023).
II.2.4 Node-line Phonons
@@
Phonon topological gapless states mentioned previously feature band crossings at discrete zero-dimensional points in momentum space. In certain instances, these band crossings may transform into one-dimensional curves within the three-dimensional Brillouin zone, referred to as nodal lines or nodal rings Burkov et al. (2011); Fang et al. (2016); Zhang et al. (2019b). Nodal lines or rings typically require the protection of additional symmetries, such as mirror symmetry, product of spatial inversion and time-reversal symmetry ($\mathcal{PT}$), and the like. Perturbations that conserve these symmetries cannot eliminate the crossing line and thereby cannot open a complete energy gap. Consider a two-band phonon system that preserves mirror symmetry $M_{z}$; the low-energy effective Hamiltonian for this system can be expressed as:
@@
In addition to the mirror-protected node-line/ring phonons, which are fixed to the mirror-invariant plane, the combined spatial inversion and time-reversal symmetry ($\mathcal{PT}$) can also safeguard nodal lines/rings. These nodal lines/rings may exist in various configurations and at any momentum within the Brillouin zone Zhang et al. (2019b). Independent of the specific symmetries of the system, node-line/ring phonons with $\gamma=\pi$ host topological surface states. These correspond to the drumhead surface states that link the nodal line/ring band crossings on the surface BZ, as depicted by the blue curve in Fig. 11 (c).
@@
II.3 Diagnosing Topological Phonons at Generic Momenta
Figure 12: The nontrivial symmetry-based indicator groups, denoted as $X_{BS}$, apply to phonon systems that preserve time-reversal symmetry. $X_{BS}$ represents the quotient group formed by the set of “gapped” band structures and the set of “atomic insulators,” a concept originally derived from electronic systems. This notion can be extended to encompass all physical systems described by band theory. Adapted from Ref. Po et al. (2017).

### v2

@@ II.3.3 Unconventional Dirac Phonons
As discussed in the previous subsection, surface arcs arise when a pair of Weyl phonons with opposite Chern numbers project to distinct momenta in the surface Brillouin zone (BZ), as illustrated in Figs. 5 (a1)-(a3). Conversely, if two Weyl points with opposite chirality project to the same momentum, the surface arcs vanish, as shown in Fig. 5 (a4). When such a pair of Weyl phonons carrying opposite Chern numbers approaches each other under $\mathcal{PT}$ symmetry, they coalesce into a Dirac node accompanied by two counter-propagating helicoidal surface states. Crucially, this $\mathcal{PT}$-protected Dirac phonon carries a net monopole charge of zero, rendering it topologically trivial and devoid of protected surface states in Dirac semimetals kargarian2016surface ; le2018dirac . Introducing a symmetry-preserving perturbation hybridizes the anti-parallel surface states, lifting their degeneracy and generating gapped, topologically trivial surface states kargarian2016surface ; le2018dirac . Despite this straightforward theoretical picture, two critical challenges remain unresolved before the year of 2018:
@@
(p1) In electronic systems, double Fermi arcs have predominantly been proposed in Dirac semimetals like Na3Bi wang2012dirac ; xiong2015evidence ; liu2014discovery and Cd3As2, wang2013three ; liu2014stable ; neupane2014observation ; yi2014evidence ; he2014quantum ; borisenko2014experimental ; li2016negative , they serve as critical experimental signatures of their topological nature. These arcs arise from the hybridization of two anti-parallel surface states, which remains incomplete due to their weak hybridization. A key unresolved question is identifying the conditions that lead to fully gapped surface states (and the absence of Fermi arcs). In other words, what is the microscopic mechanism for obtaining a strong hybridization of two anti-parallel surface states in Dirac phonon systems?
@@
To address (p1), le2018dirac pointed out that the deformation of the anti-parallel surface states (termed double Fermi arcs in their study) arises due to a significant $k^{3}$ term that respects all crystal symmetries. They illustrated this with first-principle calculations on $\beta$-CuI, for which the effective Hamiltonian at the Dirac point can be formulated as $H_{\text{eff}}(\bm{k})$:
@@
To address (p2), a new type of topological invariant termed $\mathbb{Z}_{2}$-type monopole charge $Q$ must be introduced for topological Dirac phonons, since the $\mathbb{Z}$-type monopole charge $C$ is zero. fang2016topological ; zhang2022z ; zhang2023parallel ; Yukitake2024redefinition have highlighted that in systems exhibiting $\mathcal{TG}$ symmetry, where $\mathcal{G}$ signifies the glide mirror symmetry, this $\mathbb{Z}_{2}$ monopole charge $Q$ can be defined fang2016topological ; cheng2020discovering ; zhang2022z ; zhang2023parallel ; Yukitake2024redefinition ; su2022three ; cai2020symmetry ; qian2023stable ; lu2016symmetry ; kim2021theoretical ; yukitake2025double . When $Q\neq 0$, double helicoid surface states emerge, as shown in Fig. 8 (a3). Conversely, if $Q=0$, the Dirac phonon is trivial with gapped surface states. Details are as follows.
@@
Figure 7: The $\mathbb{Z}_{2}$ monopole charge $Q$ is defined using specific surfaces and lines. In (a), $Q$ is defined on a surface $S$ that does not intersect with the plane $k_{z}=0$. This surface $S$ is closed under $\hat{\Theta}$, with the half part $S_{+}$ ($k_{z}\geq\pi$) shown in (b). The boundary of $S_{+}$, denoted as $\partial S_{+}=l_{1}\cup l_{2}$, consists of two lines. Examples of such surfaces $S$ are provided in (c) and (d), where $S$ is defined by a torus $T$ and a sphere $S^{2}$, respectively. (e) illustrates a closed line $l$ used to define the quantity $P[l]$. Adapted from Ref. Yukitake2024redefinition .
@@
Yukitake2024redefinition considered a system that possess a single glide symmetry operation, denoted as $\mathcal{G}_{y}=\{M_{y}|00\frac{1}{2}\}$, as shown in Fig. 8 (a). When the time-reversal symmetry, $\mathcal{T}$, is a symmetry of the system, the operation $\tilde{\Theta}_{y}$ $\mathcal{T}$$\mathcal{G}_{y}$ is also conserved. It is found that $\tilde{\Theta}_{y}^{2}=e^{-ik_{z}}$, indicating that the $k_{z}=\pi$ ($\pi$) plane is distinctive, as it satisfies the condition $\tilde{\Theta}_{y}^{2}=-1$. In this context, Yukitake2024redefinition examined the $\mathbb{Z}_{2}$-type Dirac points on the $k_{z}=\pi$ plane in the 3D BZ. These two Dirac points are connected by $\mathcal{T}$} symmetry, precluding their occurrence at time-reversal-invariant momenta (TRIM). Consequently, $\mathbb{Z}_{2}$ Dirac points typically emerge along high-symmetry lines in systems possessing a single glide mirror symmetry, and at non-TRIM high-symmetry points in systems with two vertical glide mirror symmetries fang2016topological ; zhang2022z ; zhang2023parallel ; Yukitake2024redefinition ; kim2021theoretical ; yukitake2025double . Subsequently, we will review the convention-independent $\mathbb{Z}_{2}$ monopole charge $QS$ associated with an oriented closed surface $S$, which adheres to the following criteria as established by Yukitake2024redefinition : (i) The surface $S$ is closed under the operation $\hat{\Theta}:(k_{x},k_{y},k_{z})\mapsto(-k_{x},k_{y},-k_{z})$, (ii) The surface $S$ does not intersect with the plane defined by $k_{z}=0$ (iii) The system exhibits a bulk energy gap over the entire surface $S$.
@@
Analogous to $\mathbb{Z}$-type monopole charge C{C}, the $\mathbb{Z}_{2}$-type monopole charge Q{Q} is defined in terms of wavefunctions over a sphere $S$ that encloses the Dirac point, as described in fang2016topological . Thus, Q{Q} can be expressed as:
$Q=\frac{1}{2\pi}\int_{S_{+}}d\bm{S}\cdot\nabla_{\bm{k}}\times\bm{A}(\bm{k})-2\sum_{l\in\partial S_{+}}P[l]\quad(\operatorname{mod}2),$ (42)
@@
$(-1)^{Q}=\prod_{l\in\partial S_{n}}\frac{\operatorname{Pf}{\omega}[l(0)]}{\sqrt{\det\omega[l(0)]}}\frac{\operatorname{Pf}{\omega}[l(\pi)]}{\sqrt{\det\omega[l(\pi)]}}.\$(43)
@@
Thus, $Q[T]$ is the difference of $\tilde{\Theta}-$polarization, and it serves to establish the bulk-surface correspondence of $\mathbb{Z}_{2}$ Dirac points by considering the surface $T$ as a 2D BZ $\pi,\pi]\times[-\pi,\pi]$ in the $(s,t)$ parameter space, as shown in Fig. 8 (a1). Given that $\tilde{\Theta}^{2}=-1$ along the lines where $t=0$ and $t=\pi$, the anti-unitary operator $\tilde{\Theta}$ behaves analogously to time-reversal symmetry in $\mathcal{T}$-symmetric $\mathbb{Z}_{2}$ topological insulators fu2007topological . This implies that the energy bands, both from the bulk and the surface, will exhibit Kramers’s degeneracy at the points at $t=0$ and $t=\pi$. Upon projection onto the (100) surface BZ, the torus $T$ is projected onto a circle path $c$, as marked by the orange circle in Fig. 8 (a1). If the $\mathbb{Z}_{2}$ charge $Q[T]$ is nontrivial, an odd number of $\mathbb{Z}_{2}$ Dirac points are enclosed within the torus. Consequently, a pair of helical surface states will emerge on the circle $C$, as shown by the red and blue bands in Fig. 8 (a2). These two surface states are degenerate at $t=0$, resulting in a degenerate line along the boundary of the surface BZ. Adjusting the radius of circle $C$ allows for the emergence of two helical surface states, as shonw in Fig. 8 (a3). The intersection of these two surface states is protected by the symmetry $\mathcal{GT}$, leading to the formation of double helical (anti-parallel) surface states. Figure 8 (a4) illustrates the Fermi surface of a system hosting two $\mathbb{Z}_{2}$ Dirac points, featuring two Fermi arcs that link the projection of the Dirac points.
@@
$\mathbb{Z}_{2}$-type Dirac points may reside at various momenta, subject only to the constraint of satisfying the $\tilde{\Theta}_{y}$ symmetry, which is associated with a high-symmetry line in the BZ. When two Dirac points are projected onto distinct momenta within the surface BZ, a pair of anti-parallel helical surface states emerges. These states are topologically protected by the charge Q{Q}, as shown in Fig. 8 (b2). When two $\tilde{\Theta}_{i,j}$ symmetries are present, with $i,j\in\{x,y,z\}$, Dirac points are confined to the high-symmetry points within the BZ, as listed in Tab. 8. In this case, if two Dirac points are projected onto the identical momentum within the surface BZ, the quadruple surface states appear, as shown in Fig. 8 (b3).
@@
In certain instances, a $\mathbb{Z}_{2}$ Dirac point may split into a pair of Weyl dipoles or transform into a $\mathbb{Z}_{2}$ nodal ring, a transformation that depends on the crystalline symmetries inherent to the system. All of the topological band crossings mentioned here can be cataloged by the $\mathbb{Z}_{2}$-type monopole charge Q{Q}, associated with topologically protected surface states fang2016topological ; cheng2020discovering ; zhang2022z ; zhang2023parallel ; Yukitake2024redefinition ; su2022three ; cai2020symmetry ; qian2023stable ; kim2021theoretical ; yukitake2025double ; lu2016symmetry .
@@
Figure 8: “Bulk-surface correspondence” for the $\mathbb{Z}_{2}$ monopole charge $Q[T]$, defined on the torus $T$. (a1) Bulk and (100) surface Brillouin Zones. A $\mathbb{Z}_{2}$ Dirac point is situated within the $k_{z}=\pi$ plane of the torus $T$ and is projected onto the circle $\mathcal{C}$. The $k_{z}=\pi$ plane is characterized by $\mathcal{TG}^{2}=-1$. The torus may be conceptualized as a 2D rillouin Zone in the parameter space of ($s,t$), where Krammers’ degeneracy appears at $t=\pi,0$. (a2) When a $\mathbb{Z}_{2}$ Dirac point is enclosed within $T$, the surface states along the loop $\mathcal{C}$ exhibit helical characteristics and are degenerate at $t=0$ ($t=\pi$). (a3) By altering the radius of the loop $\mathcal{C}$, double-helicolid surface states is obtained, associated with two Fermi arcs shown in (a4). The charge $Q[T]$ is applicable to systems with $\mathbb{Z}_{2}$-charged nodal ring or a Weyl dipole, which can interconvert with Dirac points. (b1) Represents the topological surface state connecting a pair of Weyl phonons with opposite chirality. (b2) and (b3) are the double and quadruple helical surface states, respectively, which are attributed to the $\mathbb{Z}_{2}$ Dirac phonons. (a) is adapted from Ref. Yukitake2024redefinition and (b) is adapted from Ref. zhang2023parallel .
II.3.4 Node-line Phonons
@@
Phonon topological gapless states mentioned previously feature band crossings at discrete zero-dimensional points in momentum space. In certain instances, these band crossings may transform into one-dimensional curves within the three-dimensional Brillouin zone, referred to as nodal lines or nodal rings burkov2011topological ; fang2016topological ; zhang2019phononic ; li2025general . Nodal lines or rings typically require the protection of additional symmetries, such as mirror symmetry, product of spatial inversion and time-reversal symmetry ($\mathcal{PT}$), and the like. Perturbations that conserve these symmetries cannot eliminate the crossing line and thereby cannot open a complete energy gap. Consider a two-band phonon system that preserves mirror symmetry $M_{z}$; the low-energy effective Hamiltonian for this system can be expressed as:
@@
In addition to the mirror-protected node-line/ring phonons, which are fixed to the mirror-invariant plane, the combined spatial inversion and time-reversal symmetry ($\mathcal{PT}$) can also safeguard nodal lines/rings. These nodal lines/rings may exist in various configurations and at any momentum within the Brillouin zone zhang2019phononic . Independent of the specific symmetries of the system, node-line/ring phonons with $\gamma=\pi$ host topological surface states. These correspond to the drumhead surface states that link the nodal line/ring band crossings on the surface BZ, as depicted by the blue curve in Fig. 9 (c).
@@
II.4 Diagnosing Topological Phonons at Generic Momenta
Table 4: The nontrivial symmetry-based indicator groups, denoted as $X_{BS}$, apply to phonon systems that preserve time-reversal symmetry. $X_{BS}$ represents the quotient group formed by the set of “gapped” band structures and the set of “atomic insulators,” a concept originally derived from electronic systems. This notion can be extended to encompass all physical systems described by band theory. Adapted from Ref. po2017symmetry .
$X_{BS}$ Space groups
$\mathbb{Z}_{2}$ 3, 11, 14, 27, 37, 48, 49, 50, 52, 53, 54, 56, 58, 60, 66, 68, 70, 75, 77, 82, 85, 86, 88, 103, 124, 128, 130, 162, 163, 164, 165, 166, 167, 168, 171, 172, 176, 184, 192, 201, 203
$\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 12, 13, 15, 81, 84, 87
$\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ 147, 148
$\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}$ 10, 83, 175
$\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{2}\times\mathbb{Z}_{4}$ 2